# Berezinskii-Kosterlitz-Thouless transition in the XY model on the honeycomb lattice: A comprehensive Monte Carlo analysis

### v3

@@ V Conclusions
We carefully analyzed the XY model on the honeycomb lattice using the numerical Monte Carlo method, applying two techniques: the Simulated Annealing protocol and entropic simulations based on the Wang-Landau algorithm. Our primary objective was to calculate the critical transition temperature and to compare our findings with theoretical predictions. To achieve this, we used the second- and fourth-order helicity modulus as order parameters, as these properties exhibit a discontinuity during the phase transition. Our best estimates were obtained using the SA protocol, which provided values for the inverse transition temperature of $\beta_{BKT}=1.635(11)$ (based on fourth-order helicity) and $T_{BKT}=1.687(3)$ (based on second-order helicity). Notably, these values are significantly higher than the analytical prediction of $\sqrt{2}$. By analyzing the crossing of the Binder cumulant, we obtained an even higher value of $\beta_{cross}=1.724\pm 0.002$. This value is closer to the results obtained through neural network calculations. Additionally, we calculated the vortex density, which allowed us to estimate the formation energy of a pair of vortices. The obtained value was $2\mu=4.0\pm 0.5$, in units of $J$. This is more than 50% smaller than that of the square lattice.
Our findings indicate that theoretical calculations of the critical temperature for the XY model on the honeycomb lattice need to be revisited. Also, more precise numerical calculations are necessary to resolve the inconsistencies among different techniques.

### v4

@@ V Conclusions
We conducted a thorough analysis of the XY model on a honeycomb lattice using the numerical Monte Carlo method, employing two techniques: the Simulated Annealing protocol and high-resolution entropic simulations based on the Wang-Landau algorithm. Our primary goal was to calculate the critical transition temperature and to compare our results with theoretical predictions.
To achieve this, we utilized the second- and fourth-order helicity modulus as order parameters, as these properties exhibit a discontinuity at the BKT phase transition. Our best estimates for the transition temperature were obtained using the helicity modulus from both techniques: we found $T_{BKT}=0.575(8)$ from the Simulated Annealing method and $T_{BKT}=0.576(3)$ from the Wang-Landau algorithm.
Furthermore, we assessed the applicability of the fourth-order helicity modulus in determining the BKT temperature. While it presents a minimum at the transition that scales with system size, the large fluctuations in the data obtained from Simulated Annealing hinder the precise identification of the critical temperature. In contrast, the data from the Wang-Landau method is more precise and smoother, although it is limited to larger system sizes, yielding values of $T_{BKT}=0.551(11)$ from Simulated Annealing and $T_{BKT}=0.568(1)$ from Wang-Landau.
Interestingly, these results align closely with those obtained independently through neural network calculations. Notably, these values are significantly lower than the analytical prediction of $1/\sqrt{2}$. Our findings suggest that theoretical calculations of the critical temperature for the XY model on the honeycomb lattice need to be reevaluated.
Additionally, we calculated the vortex density, which allowed us to estimate the formation energy of a pair of vortices. The obtained value was $2\mu=5.80(12)$ in units of $J$, which is smaller than the corresponding value for the square lattice, justified by the idea of instability of the honeycomb lattice to support the spin long-range order.